INAF queue observing, 2017 Feb 18-19

Observer: R. Carini, F. Cusano
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: D. Thompson on call
Instrument:

Summary

Snow Showers and Breezy

The dome stays closed.

Night duration (12\xB0 el twilight) 02:30am-05:45pm UT = 10 h 15m

Observing time: 0 hrs 0 m (0%)

Weather time loss: 10 hrs 15m (100%)

technical time loss 0 m (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
LUCI CALIBRATION (DARK)        

Details

01:50 LUCI1 calibrations DARK
  • luci.20170219.0001-0005 dit=300, ndit=1, mer, norm. NO GOOD
  • luci.0006-0010 dit=300, ndit=1, mer, norm
  • luci.0011-0015 dit=5, ndit=6, lir, int
  • luci.0016-0020 dit=4, ndit=2, lir, int
  • luci.0021-0025 dit=240, ndit=1, mer, norm
  • luci.0026-0030 dit=270, ndit=1, mer, norm
  • luci.0031-0035 dit=200, ndit=1, mer, norm
  • luci.0035-0040 dit=150, ndit=1, mer, norm
  • luci.0041-0045 dit=120, ndit=1, mer, norm
  • luci.0046-0050 dit=170, ndit=1, mer, norm
  • luci.0051-0055 dit=220, ndit=1, mer, norm
  • luci.0056-0060 dit=180, ndit=1, mer, norm
  • luci.0060-0065 dit=220, ndit=1, mer, norm

FLAT * error while trying to put the mask in FPU * we can not use LUCI1 during the night

-- %USERSIG{RobertaCarini - 2017-02-18}%

Comments

 
Topic revision: r4 - 21 Feb 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback